1.86 cm3 of a liquid of density 0.803 g cm-3was injected into a syringe maintained at 200°C.
The liquid vapourised completely and occupied 808.54 cm3 at 1.20 atm.

a) Calculate the volume that this sample of vapour will occupy at STP

b) Hence calculate the molar mass of the liquid
1
Expert's answer
2018-03-06T07:42:01-0500
a) M=1.86*0.803=1.49358 g

V(STP)=(273/473)*(1.2/1.0)*808.54=560.0 cm-3

N=V/Vm=560/22400=0.025 mol

b) M=m/n=1.49358:0.025=59.7